Compared to Tunisia, Malaysia's GDP per capita is USD9945 for 2017. What is the average growth rate that Tunisia should maintain since 2009, if the country wants to achieve GDP per capita in line with Malaysia?
1
Expert's answer
2022-02-17T10:02:36-0500

Tunisia GDP= 3687.77

Difference Malaysia= 9945- 3687.77= 6257.77

"Growth rate= \frac{6257.77}{3687.77}\times 100= 169.7"